OpenStudy (anonymous):

find the value of x in the exponential equation 2^[(2/3)x+1] – 3 . 2^[(1/3)x] – 20 = 0 I'm having a hard time here.. can somebody help me??

OpenStudy (ikram002p):

the equation is like this right ?? \[\left( 2^\frac{ 2x+3 }{ 3 }-3 \right)\left( 2^\frac{ x }{3 } -30\right)=0\]

OpenStudy (ikram002p):

-20*

OpenStudy (anonymous):

thanks:D

OpenStudy (ikram002p):

that was not the answer ! I was cheking from the question

OpenStudy (anonymous):

im just thanking because ive waited hours before somebody entertains my question

OpenStudy (ikram002p):

ok lets do it together ,first lets analyze brackets \[2^\frac{ x-3 }{ 3 } -(2^\frac{ 2x-1 }{ 3} \times 5)-(2^\frac{ x }{ 3 }\times3)+60\]

OpenStudy (anonymous):

did you expand it? im bit confused

OpenStudy (ikram002p):

yeb , I did , whats make u confuse ?

OpenStudy (ikram002p):

mmm Im going to sol the question in different way

OpenStudy (ikram002p):

\[\left( 2^\frac{ 2x-3 }{ 3 } -3\right)\left( 2^\frac{ x }{ 3 } -20\right)=0\]

OpenStudy (ikram002p):

assume \[y=2^\frac{ x }{ 3 }\] \[y^2=2^\frac{ 2x }{ 3 }\] the equation convert to \[(2y^2-3)(y-20)=0\] case one \[y=20 \] when y=20 \[2^\frac{ x }{ 3 }=20\] case two \[y^2=\frac{ 3 }{ 2 } ,2^\frac{ 2x }{ 3 }=\frac{ 3 }{ 2 }\]

OpenStudy (anonymous):

\[the equation is like this 2^\left( 2/3)(x+1 \right)-3\times2^\left( 1/3)(x \right)-20=0\]

OpenStudy (ikram002p):

now its better

OpenStudy (ikram002p):

like this ? \[2^\frac{ 2x+2 }{ 3 }-3\times2^\frac{ x }{ 3 }-20=0\] or this ? \[2^\frac{ 2x+3 }{ 3 }-3\times2^\frac{ x }{ 3 }-20=0\]

OpenStudy (anonymous):

the first one

OpenStudy (ikram002p):

ok assume \[y=2^\frac{ x }{3 }\] \[y^2=\frac{ 2x }{ 3 }\] the equation would be \[2^\frac{ 2 }{ 3 }\times y^2-3y-20=0\] solve for y using the general law

OpenStudy (anonymous):

where did the y came from? i thought we're finding for x?

OpenStudy (ikram002p):

I assumed \[y=2^\frac{ x }{ 3}\]

OpenStudy (anonymous):

oh ok..

OpenStudy (ikram002p):

so finally y=6 or 0 when y=6 \[2^\frac{ x }{ 3 }=6\] or \[2^\frac{ x }{ 3 }=0\] is that ok untill now ??

OpenStudy (anonymous):

yes.. keep going

OpenStudy (ikram002p):

\[x=\ln \sqrt[3]{6},for.base2\rightarrow or, x=\ln 0, for.base 2\]

OpenStudy (anonymous):

uhmm i remember the process thanks for refreshing it ... im very glad that you lend me your time.. i better get going because my mom just scolded me earlier.. i hope you'll teach me again next time :D

OpenStudy (ikram002p):

any thing u need I'd be happy if I help u :)